HUBBLE SPACE TELESCOPE DAILY REPORT #1405 PERIOD COVERED: 0000Z (UTC) 06/05/95 - 0000Z (UTC) 06/06/95 Daily Status Report as of 157/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Four Sets of WF/PC-2 5562 (WF/PC-2 Cycle 4 Calibration: Internal Monitor 3: Darks) The WF/PC-2 was used to obtain several 30-minute long, dark exposures to provide ongoing calibration of the CCD dark current and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 4819 (Point Spread Function Characterization) The WF/PC-2 was used to make multiple exposures of the open cluster M-67 in Cancer (aka NGC-2682) during coarse and fine dwell scans to study the point spread function. A dark calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed FOS/BL 6146 (FOS Location of Spectra and Aperture - Part 2) The Faint Object Spectrograph (Blue Side) was used to conduct a test to determine the repeatability of positioning the filter-grating wheel. All observations were internal. A wave length calibration and a dark image were also taken. All observations were executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 5980 (The Jet of 3C273) The Planetary Camera was used to image the jet of 3C273. The objective is to study the morphology of the jet at high resolution. The investigators are particularly interested in comparing these data with optical and radio maps to investigate the mechanism for particle acceleration, which occurs all along the jet. The proposal completed with no reported anomalies. 1.5 Completed WF/PC-2 5505 (Deep Probes of Gravitational Lens Systems) The WF/PC-2 was used to make deep images of of the elliptical galaxy MG11654+134 which is a member of the Coma Cluster. The intent is to improve our understanding of the stellar populations in ellipticals and to determine the distance to the Coma Cluster. Although the acquisition failure to coarse track (described in section 2.1) occurred during this proposal, the observations were completed as planned, and no problems were noted. 1.6 Completed WF/PC-2 5655 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 2: Flats Part 2) The WF/PC-2 was used to make both dark and visual flats. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 At 156/1235Z, an acquistion failed to Coarse Track on Tracker 3 when presence was not achieved on Tracker 2. The SMS magnitude was 9.498, and the estimated magnitude was 9.61. The tracker remained in coarse track until 157/0038Z, with occultations. HSTAR 5215 was written to document this event. Scheduled Reacquisitions : 11 Successful : 11 2.2 FHST Updates: Scheduled : 9 Successful : 8 One failure occurred at 157/0828Z. THe reason is currently unknown, and analysis is ongoing. 2.3 Significant HSTARs: None 2.4 Operations Notes: No significant operational problems were reported. 3.0 SIGNIFICANT FORTHCOMING EVENTS: The RMGA will be powered on and off today for the interval 157/1200Z - 1521Z. /DMH -END-